OpenStudy (anonymous):

domain of square root (5x-20) in interval notation

OpenStudy (anonymous):

First, can you figure out the domain of \[\sqrt{x}\]?

OpenStudy (anonymous):

Think about positive values of X versus negative values of X.

OpenStudy (anonymous):

You can take the square root of a positive number, or 0, but you can't take the square root a negative number. That means \[x \ge 0\] or in interval notation, since this includes 0 and goes upwards forever, is \[[0, \infty)\]

OpenStudy (anonymous):

All of that means that what's inside of the square root in your question has to be greater than or equal to 0. \[5x-20 \ge 0\]

OpenStudy (anonymous):

You should be able to figure out the rest from there pretty easily.